In the circuit shown, R1 = 7 ohm, R2 = 11 ohm, R3
Answer #1

Here ,

the current in resistor R1 is I1

current in resistor is R2 is I2 ,

using KVL

19 - 7 * I1 - 9 *(I1 - I2) = 0

16 I1 - 9 I2 = 19 ----(1)

again using KVL

- 11 * I2 + 8 - 9 (I2 - I1) = 0

-9 I1 + 20 I2 = 8    ----(2)

solving 1 and 2

I1 = 1.89 A

I2 = 1.25 A

Now,

V(R3) = (1.89 - 1.25) * 9

V(R3) = 5.76 V

voltage across R3 is 5.76 V